# Goal Inference from Open-Ended Dialog

### v1

@@ Abstract
We present an online method for embodied agents to learn and accomplish diverse user goals. While offline methods like RLHF can represent various goals but require large datasets, our approach achieves similar flexibility with online efficiency. We extract natural language goal representations from conversations with Large Language Models (LLMs). We prompt an LLM to role play as a human with different goals and use the corresponding likelihoods to run Bayesian inference over potential goals. As a result, our method can represent uncertainty over complex goals based on unrestricted dialog. We evaluate our method in grocery shopping and home robot assistance domains using a text-based interface and AI2Thor simulation respectively. Results show our method outperforms ablation baselines that lack either explicit goal representation or probabilistic inference.
I Introduction
@@
AI agents and robots must quickly learn and carry out many different user tasks in real-time. For example, a home robot assistant may need to adapt to various household preferences and routines. Imagine a scenario where your robot assistant is tasked with gathering ingredients to bake a cake for you. Depending on who you are, you may want different ingredients. If you only want a basic cake, your ideal recipe is eggs, milk, sugar, and flour. However, if you are allergic to gluten, you would want gluten-free flour. If you prefer strawberry cake, then you would want to get strawberries. Every human has a different set of preferences. This level of variation is challenging, if not impossible, for system designers to anticipate in advance. In order for agents/robots to perform or assist with tasks for humans, they must first be able to learn the preferences of the humans that they are trying to help.
@@
To address this challenge, we propose a new method GOOD (Goals fOr Open-ended Dialogue) that combines the best parts of offline and online approaches. It uses Large Language Models (LLMs) to infer natural language representations of user goals. This allows our method to represent a flexible, open-ended set of possible goals during an online interaction. As a result, the method combines the flexibility and representation power of offline preference tuning methods [1, 2] with the data efficiency and uncertainty quantification of online methods that learn rewards based on a set of engineered features [3, 4]. This allows our method to represent uncertainty over goals that may not have been explicitly engineered or anticipated in advance. In order to learn goals efficiently, we use natural language dialog with the user instead of, e.g., best-of-k comparisons present in [5, 6, 7, 8].
@@
However, applying traditional Bayesian inference to natural language goals presents two significant challenges. First, it is intractable to enumerate the space of all possible natural language expressions. To address this, we run Bayesian inference over a reduced set of explicit hypothesized goals. We use LLM modules to maintain this set of goals. Second, running Bayesian inference requires a likelihood function to represent the conditional distribution of a dialog for different candidate goals. Our key insight is that we can leverage an LLM’s ability to role-play as a human with an explicit goal to define this likelihood.
In this paper, we make the following contributions:
@@
II Background and Related Work
II-A Preference Learning and NL Probabilistic Reasoning with LLMs
Offline preference tuning methods [1, 2] are data heavy but are generalizable to many domains and tasks. Online methods are data efficient but often task specific [3]. Our method combines both aspects of being data efficient, can be generalized across many domains and tasks, and is online. Previous work [9, 10], show improvement with using LLMs for learning human preferences and NL probabilistic reasoning. Previous methods such as [11, 8] often rely on asking the most informative questions. Our method does not make these unrealistic assumptions on interactions and is flexible to other interaction methods. Unlike existing preference learning techniques that expect structured forms of data like yes/no queries or comparisons ([5, 6, 7, 8]), our method allows for representing goals that the designer did not explicitly engineer in advance.
@@ II-A Preference Learning and NL Probabilistic Reasoning with LLMs
Previous works such as [12, 13, 14] show that LLM agents can have different roles to achieve tasks. Our work leverages this concept and has different LLM calls that focus on specific tasks in our pipeline to be more efficient with how much information each LLM call has in memory.
II-B Human Preferences or Human Interaction with LLMs in Robotics
@@
Many works incorporate human preferences in doing robotic tasks or generating robotic plans. However, most of these either rely on best-of-k comparison ([7]), or learns rewards based on a specific set of engineered features, making it difficult to generalize to other assistance tasks or scenarios ([15]). Some works involve human-robot interactions that are much more limited or not through active conversations ([16, 17, 4]). Our work leverages natural language goals and LLMs for flexibility and understanding human preferences through language interactions.
III Method
Figure 2: Method Overview Diagram.Method Overview Diagram. The pipeline consists of four main modules: Conversation, Inference, Goal Management, and Action. The process iterates until the task is completed or the conversation limit is reached.
@@ III Method
The key idea behind our method is tracking possible human goals and how likely they are as the dialog between the human and the agent goes on. To track the possible space of human goals, we instantiate a finite goal set to which we can add goals if our inference finds that the human preferences in the conversation so far are not represented in the goals, or remove unlikely goals. Given the updated goal set, we infer the likeliest goals and select actions based on them. We continue the conversation rounds until the task is completed.
III-A Preliminaries
@@
This formulation presents two challenges for open-ended goal inference. First, how should we flexibly represent the goals themselves? Typical methods ([18], [19]) define goals as $x,y$ locations in navigation or continuous parameters trading off features in a reward function, but these approaches restrict the set of goals or preferences that can be learned.
Algorithm 1 Our Method, where $G$ is the goal list, $q$ is the robot query, $u$ is human utterance, $hp$ is the human profile, $s$ is the task status, $t$ is the transcript so far.
@@
2:while task not complete & convo num $<20$: ## Generate a round of convo $q,u,t$ = Conversation($rt,s,t,hp$) likelihoods = Belief Update(llm, $G,q,u,t$) if argmax[likelihoods] == ’Unspecified’: add($G$) remove($G$) ## update the likelihoods for new set of goals likelihoods = Belief Update(llm, $G,q,u,t$) likely goals = sort(likelihoods)[0:$k$] ## Take actions based on $k$ likely goals action history, complete = Action(llm, likely goals) if complete == True then end else: convo num +=1+=1+ = 1
Next, what should the likelihood $P(u\mid g)$ even be? Prior work has modeled this as the Boltzmann noisily rational model that assumes humans select inputs in proportion to their exponentiated reward ([20], [21]). However, this model is an oversimplification compared to reality because humans can be biased, myopic, or not even be aware of their internal reward ([22], [23]). Finally, how can we keep track of the set of goals $G$ in the denominator? Prior work has system designers define a set of possible goals ahead of time, but thinking of all possible goals in any environment or for any human preference a priori is unreasonable.
Our method consists of four modules. The Conversation Module produces a round of robot query and a human utterance. The Inference Module does inference on the goal list to find the most likely goals. The Goal Management Module is responsible for managing the goal list. The Action Module generates and takes a sequence of actions based on the most likely goals, and checks to see if the task is completed. The overview of the pipeline is shown in 2 and in Pseudocode 1. Our method elegantly tackles all these questions by leveraging natural language and powerful LLM priors.
III-B Conversation Module
We use a LLM to talk to the user about the task. The conversation module enables dialogue between a robot that asks questions (a “robot query”) and for a human that answers those questions (a “human utterance”). To generate a robot query, an LLM is prompted to generate a question given a description of the robot/agent task, transcript of the chat so far, and current status of a task. Our experiments use LLMs to roleplay as a human with a certain “human profile”. To generate a human utterance, an LLM is prompted to generate a response given the robot’s task, a human profile, the current task status, and phrase for the human to respond with if the task is completed. See Section VII for prompt details. The framework is flexible and can be easily adjusted to allow for user text input or other text generation methods.
III-C Inference Module for Belief Space Update
@@
In order to track a Bayesian posterior over a set of goals, we need to calculate the probability of a goal given a human utterance, $P(g|u)$. Using LLMs to calculate $P(g|u)$ according to Equation 1, we need a model for $P(u|g)$ – a model that tells us what utterances a person would choose conditioned on having a particular goal. LLMs are good at roleplaying ([24]) and have been demonstrated to possess strong common sense priors ([25], [26]). We leverage these strong priors to model $P(u\mid g)$ as an LLM query $\pi_{\text{LLM}}(u,g)$. By prompting the LLM to generate utterances given a goal, we can use the model’s output probabilities as an approximation of this likelihood.
@@
Consider maintaining a belief over two possible goals: “I want cocoa” and ‘I want noodles” given the utterance “I’d like something sweet.” Our method queries an LLM for the likelihood of the utterance based on the prompt ‘Role-play as a person who wants $g$’, where $g$ is replaced with each possible goal.” Given a human utterance is “I want a cake”, the likelihood of “I want cocoa” would be higher than “I want noodles”. The sum of the logits for “I want a cocoa” when the LLM is provided that “I want a cake” is the true goal is higher than the sum of the logits for “I want a cake” when the LLM is provided that “I want noodles” is the true goal. See Section VII for LLM query details.
III-D Goal Management Module
@@
To implement this module, we need to be able to propose new goals and remove unlikely goals. To propose new goals, a LLM is prompted to return a list of goals that can be added given the current goal list, the conversation transcript, and the robot’s task. We remove unlikely and unsafe goals through two ways. 1) A LLM is prompted to return a list of goals to remove that should not be taken given the robot’s task and the conversation transcript. 2) If a goal is the least likely from our inference methods for more than $n$ rounds of conversation, then it should be removed from the goal list.
@@ III-E Action Module
The action module is responsible for two things: 1) selecting action(s) that are available and taking them given the likely goals from the inference module, and 2) checking if the task is completed. This module is domain dependent, and can be easily adapted to other domains and planners. Our implementation will be further explained in more detail in the Experiments section.
III-E1 Do Action
@@
The primary purpose of this module is for the robot/agent to take actions. We do this by prompting a language model to pick actions given the most likely goals and information about their likelihood magnitudes. If the most likely goal is “Unspecified Goal”, then none of the non-unspecified goals present accurately represent the human preference, so proceed to do “no action”. If another goal is the most likely, then the current goal list represents the human preference well, so pick the top $k$ goals from the goal list. This is a benefit of our method because the action module can take actions that are useful for multiple potential likely goals. The module maintains a record of the actions that have been taken to provide context for other modules.
The LLM prompt for Do Action includes information about the previous transcript, the list of possible actions, the most likely goal, and the next $k$ likely goals in order. The prompt will return the list of actions that should be taken in order to accomplish the goal. How the list possible actions are constructed is explained more in the Experiments section.
III-E2 Task Completion Check
@@
After a sequence of actions is taken, we need to check if the task is completed according to the human’s satisfaction indicated by their utterance this round or if a certain end action is taken. This completion check is domain and task dependent. If the task is not completed, then whole pipeline is repeated for another round of conversation.
IV Experiments

### v2

@@ Abstract
Embodied AI Agents are quickly becoming important and common tools in society. These embodied agents should be able to learn about and accomplish a wide range of user goals and preferences efficiently and robustly. Large Language Models (LLMs) are often used as they allow for opportunities for rich and open-ended dialog type interaction between the human and agent to accomplish tasks according to human preferences.
In this thesis, we argue that for embodied agents that deal with open-ended dialog during task assistance:
  1. 1.
    AI Agents should extract goals from conversations in the form of Natural Language (NL) to be better at capturing human preferences as it is intuitive for humans to communicate their preferences on tasks to agents through natural language.
  2. 2.
    AI Agents should quantify/maintain uncertainty about these goals to ensure that actions are being taken according to goals that the agent is extremely certain about.
@@
We present an online method for embodied agents to learn and accomplish diverse user goals. While offline methods like RLHF can represent various goals but require large datasets, our approach achieves similar flexibility with online efficiency. We extract natural language goal representations from conversations with Large Language Models (LLMs). We prompt an LLM to role play as a human with different goals and use the corresponding likelihoods to run Bayesian inference over potential goals. As a result, our method can represent uncertainty over complex goals based on unrestricted dialog. We evaluate in a text-based grocery shopping domain and an AI2Thor robot simulation. We compare our method to ablation baselines that lack either explicit goal representation or probabilistic inference.
Acknowledgments
I would like to thank many people for supporting me along the way. First, I want to thank my advisor, Professor Dylan Hadfield-Menell, for being a great mentor and for providing amazing advice for how to be a good researcher. I look forward to continuing working together and learning from you, and also having more chats about our favorite sci-fi and fantasy series.
I would like to thank my collaborators: Professor Andreea Bobu for providing valuable feedback on the work featured on this thesis and proofreading of the original paper, and to Jingyi Qu for helping with running experiments for the Isolated Inference tests. I also am greatful for helpful and inspiring conversations about my work with Professor Jacob Andreas, Phillip J.K. Christoffersen, Pinar Ozisik, Belinda Zou Li, Michelle Li, Seiji Shaw, and Andi Peng.
I thank my labmates at the Algorithmic Alignment Group for their questions and feedback during group meetings, for amazing coding tool suggestions, fun lab shenanigans, and enduring through the times I dragged them down the hall to get Generative Boba.
I also thank my undergraduate research advisors, Professors Stefanie Tellex and George Konidaris, for their support during the time I was active in their labs at Brown University as an undergrad and first starting to do robotics research. I thank those from H2R and IRL labs for inspiring me to work in HRI and with language (Ifrah Idrees, Thao Nguyen, Ben Spiegel), for using AI2Thor (Eric Rosen), for showing that robots are fun but take a while to work with (Ben Abbatematteo, Max Merlin), and for helping me submit my first paper (David Paulius).
I thank those involved in the robotics communities and GW6 (Grad Women in Course 6) at MIT for helping me feel welcome and for fun times. Special shoutout for Rachel Holladay “Rachel the Elder” for being a great peer mentor, a role model, and also for all the fun at robotics conferences we’ve created when telling people our name. I also thank my friends for fun chats, lunches, tea time, karaoke nights, movie nights, coffee hours, and chaotic board game nights.
Finally, I am thankful for my parents and their never ending support.
1 Introduction
@@
AI agents and robots must quickly learn and carry out many different user tasks in real-time. Imagine a scenario where your robot assistant is tasked with gathering ingredients to bake a cake for you. Depending on who you are, you may want different ingredients. If you only want a basic cake, your ideal recipe is eggs, milk, sugar, and flour. However, if you are allergic to gluten, you would want gluten-free flour. If you prefer strawberry cake, then you would want to get strawberries. Every human has a different set of preferences. This level of variation is challenging, if not impossible, for system designers to anticipate in advance. In order for agents/robots to perform or assist with tasks for humans, they must a) be able to learn the preferences of the humans that they are trying to help and b) be able to accommodate learning these preferences over multi-rounded dialog with the humans.
In this thesis, we argue that for embodied agents that deal with open-ended dialog during task assistance:
  1. 1.
    AI Agents should extract goals from conversations in the form of Natural Language (NL) to be better at capturing human preferences as it is intuitive for humans to communicate their preferences on tasks to agents through natural language.
  2. 2.
    AI Agents should quantify/maintain uncertainty about these goals to ensure that actions are being taken according to goals that the agent is extremely certain about.
@@
We present a new method GOOD (Goals fOr Open-ended Dialogue) that combines the best parts of offline and online approaches. It uses Large Language Models (LLMs) to infer natural language representations of user goals. This allows our method to represent a flexible, open-ended set of possible goals during an online interaction. As a result, the method combines the flexibility and representation power of offline preference tuning methods Ouyang et al. (2022); Stiennon et al. (2020) with the data efficiency and uncertainty quantification of online methods that learn rewards based on a set of engineered features Biyik and Sadigh (2018); Jain et al. (2015).
This allows our method to represent uncertainty over goals that may not have been explicitly engineered or anticipated in advance. In order to learn goals efficiently, we use natural language dialog with the user instead of, e.g., best-of-k comparisons present in Rafailov et al. (2024); Yuan et al. (2023); Kuleshov and Ellis (2023); Handa et al. (2024).
@@
However, applying traditional Bayesian inference to natural language goals present two significant challenges. First, it is intractable to enumerate the space of all possible natural language expressions. To address this, we run Bayesian inference over a reduced set of explicit hypothesized goals. We use LLM modules to maintain this set of goals. Second, running Bayesian inference requires a likelihood function to represent the conditional distribution of a dialog for different candidate goals. Our key insight is that we can leverage an LLM’s ability to role-play as a human with an explicit goal to define this likelihood.
We make the following contributions:
@@
Chapter 2 presents a summary of related work in language grounding and probabilistic reasoning for agents/robots. Chapter 3 covers our method: our proposed Bayesian update equation, and the four modules: Conversation Module, Inference Module, Goal Management Module, and the Action Module. Chapter 4 provides details and discussion about implementation and evaluations for the text-based grocery experiment and AI2-Thor robot simulation experiments. Chapter 5 covers future work and conclusion.
Content of this thesis has been presented at the GenAI-HRI workshop at RSS 2024 and is under submission at another conference venue.
2 Background and Relevant Work
For robots/embodied agents that can achieve difficult tasks while interacting with humans (ie: household tasks), language grounding for robots is important – that robots can understand language and make connections between language and its physical environment. For example, if a human user wants the robot to retrieve a strawberry cake, then the robot should 1) identify the cake from the rest of the objects from its vision data 2) and then also be able to produce a robot trajectory to be able to move the strawberry cake to the human. There are also other task requirements that are not explicitly mentioned by the human but that is implied, which makes this a difficult problem: that the robot has to also be able to identify the human’s location in its vision data and other obstacles to avoid while delivering the cake, and also the actions to take to make sure that the cake arrives the right side up and intact to the human.
Before the rise of Large Language models (LLMs), Combinatory Context Grammers (CCGs) and formal semantic languages for task specification/commands highlighted in the survey paper Tellex et al. (2020) were often used, such as LTL (Linear Temporal Logic)Emerson (1990) for RL agents Toro Icarte et al. (2018); Littman et al. (2017). Many formal semantic languages do help with guarantees and safety in robot behavior Kress-Gazit et al. (2018), however they can be very unintuitive to use for a typical human user Pakonen et al. (2016), or can accommodate a limited number of expressions compared to the complete flexibility of natural language. Large databases such as WordNetFellbaum (1998), VerbNetSchuler (2005), FrameNetBaker et al. (1998), ImageNetDeng et al. (2009) were also often used for grounding language. There are also learning based methods, such as relevant object class-action pairs, segmented objects and natural language descriptions, text descriptions and 3D shapes Nguyen et al. (2020); Chen et al. (2019); Cohen et al. (2019). However, learning based methods are often domain or task specific.
Our work fits in the intersection of learning human preferences (probabilistic reasoning over dialog with LLMs) and robotic planning.
Markov Decision Process (MDP), are often used for decision making for robotics. It is represented by the tuple ($S,A,R,T$). $S$ is the state space, $A$ is the action space, $R$ is a reward function, and $T$ is a transition function. $R$ is a reward function from taking action $a$ from some transition from state $s$ to $s^{\prime}$, and $T$ is the probability function of getting to $s^{\prime}$ when taking action $a$ for state $s$. The goal of the MDP is to find a policy $\pi$ that can maximize the expected sum of rewards Sutton and Barto (2018).
@@ 2 Background and Relevant Work
The rise of LLMs for embodied agents/robots came as LLMs are realized to have extensive commonsense knowledge which can be helpful for better reasoning and decision making Zhao et al. (2024). LLMs have started to become more used for used for few-shot task planning for large environments for embodied agents Huang et al. (2022a); Ahn et al. (2022); Huang et al. (2022b). Previous works such as Park et al. (2023); Qian et al. (2023); Hong et al. (2023); Shen et al. (2024) show that LLM agents can have different roles to achieve tasks. Our work leverages this concept and has different LLM calls that focus on specific tasks in our pipeline to be more efficient with how much information each LLM call has in memory. Park et al. (2023) also show that LLMs can be used to simulate human behaviors, something we take advantage of for our experiments in our work.
@@
Wu et al. (2023); Doğan et al. (2022); Singh et al. (2023); Ren et al. (2023); Doğan et al. (2022); Biyik and Sadigh (2018); Jain et al. (2015); Kuleshov and Ellis (2023) incorporate human preferences in doing robotic tasks and generating robotic plans. However, most of these either rely on best-of-k comparison, or learns rewards based on a specific set of engineered features, making it difficult to generalize to other tasks or scenarios. Our work leverages natural language goals and LLMs for flexibility.
Initially, LLMs had trouble with reacting to varying human instructions and were sometimes even misaligned Stiennon et al. (2022); Weidinger et al. (2021); Kenton et al. (2021); Bommasani et al. (2021), so human preference learning methods are used/developed for LLMs such as Ouyang et al. (2022); Knox and Stone (2008); Bai et al. (2022); Christiano et al. (2017); Askell et al. (2021) and other methods mentioned in the survey Jiang et al. (2024).
Offline preference tuning methods Ouyang et al. (2022); Stiennon et al. (2020) are data heavy but are generalizable to many domains and tasks. Online methods are data efficient but often task specific Biyik and Sadigh (2018). Our method combines both aspects of being data efficient, can be generalized across many domains and tasks, and is online. Previous work Li et al. (2023); Austin et al. (2024), show improvement with using LLMs for learning human preferences and NL probabilistic reasoning. Previous methods such as Grand et al. (2024); Handa et al. (2024) often rely on asking the most informative questions. Our method does not make these unrealistic assumptions on interactions and is flexible to other interaction methods. Unlike existing preference learning techniques that expect structured forms of data like yes/no queries or comparisons Rafailov et al. (2024); Yuan et al. (2023); Kuleshov and Ellis (2023); Handa et al. (2024), our method allows for representing goals that the designer did not explicitly engineer in advance.
Other works that explore Bayesian inference models for cooperative teaming, goal inference, task assistance with the help of probabilistic programs include Ying et al. (2023); Zhi-Xuan et al. (2024a, b).
3 Method for Goal Inference from Open-Ended Dialog
@@
The key idea behind our method is tracking possible human goals and how likely they are as the dialog between the human and the agent goes on. To track the possible space of human goals, we instantiate a finite goal set to which we can add goals if our inference finds that the human preferences in the conversation so far are not represented in the goals, or remove unlikely goals. Given the updated goal set, we infer the most likely goals and select actions based on them. We continue the conversation rounds until the task is completed (or we’ve reached the upper conversation limit).
3.1 Preliminaries
@@
This formulation presents two challenges for open-ended goal inference. First, how should we flexibly represent the goals themselves? Typical methods Fisac et al. (2018); Bobu et al. (2020) define goals as $x,y$ locations in navigation or continuous parameters trading off features in a reward function, but these approaches restrict the set of goals or preferences that can be learned.
Next, what should the likelihood $P(u\mid g)$ even be? Prior work has modeled this as the Boltzmann noisily rational model that assumes humans select inputs in proportion to their exponentiated reward Baker et al. (2007); Jaynes (1957). However, this model is an oversimplification compared to reality because humans can be biased, myopic, or not even be aware of their internal reward Chan et al. (2021, 2019). Finally, how can we keep track of the set of goals $G$ in the denominator? Prior work has system designers define a set of possible goals ahead of time, but thinking of all possible goals in any environment or for any human preference a priori is unreasonable.
Algorithm 1 Our Method, where $G$ is the goal list, $q$ is the robot query, $u$ is human utterance, hphp is the human profile, $s$ is the task status, $t$ is the transcript so far.
@@
2:while task not complete & convo num <20<20: ## Generate a round of convo $q,u,t$ = Conversation($rt,s,t,hp$) likelihoods = Belief Update(llm, $G,q,u,t$) if argmax[likelihoods] == ‘Unspecified’ then add($G$) remove($G$) ## update the likelihoods for new set of goals likelihoods = Belief Update(llm, $G,q,u,t$) likely goals = sort(likelihoods)[0:$k$] ## Take actions based on $k$ likely goals action history, complete = Action(llm, likely goals) if complete == True then end else: convo num +=1+=1
3.2 Method
Our method consists of four modules. The Conversation Module produces a round of robot query and a human utterance. The Inference Module does inference on the goal list to find the most likely goals. The Goal Management Module is responsible for managing the goal list. The Action Module generates and takes a sequence of actions based on the most likely goals, and checks to see if the task is completed. A diagram of the overview of the pipeline is shown in Figure 5 and in Algorithm Pseudocode 1. Our method elegantly tackles all these questions by leveraging natural language and powerful LLM priors.
Conversation Module
We use a LLM to talk to the user about the task. The conversation module enables dialogue between a robot that asks questions (a “robot query”) and for a human that answers those questions (a “human utteranc”). To generate a robot query, an LLM is prompted to generate a question given a description of the robot/agent task, transcript of the chat so far, and current status of a task. Our experiments use LLMs to roleplay as a human with a certain “human profile”. To generate a human utterance, an LLM is prompted to generate a response given the robot’s task, a human profile, the current task status, and phrase for the human to respond with if the task is completed. The framework is flexible and can be easily adjusted to allow for user text input or other text generation methods. Figure 2 contains templates of the prompt we used for the LLM to generate the robot query and the human response.
Figure 2: Conversation Module LLM Prompts for LLMs to generate the robot query and the synthetic human response for the conversation module. The robot query is generated based on descriptions of the agent and the task, current status of the task, and the transcript of the conversation so far. The human response is based on the description and current status of the task, a human profile (preferences of the human), and information about the completion requirements for the task.
Inference Module for Belief Space Update
@@ 3.2 Method
In order to track a Bayesian posterior over a set of goals, we need to calculate the probability of a goal given a human utterance, $P(g|u)$. Using LLMs to calculate $P(g|u)$ according to Equation 1, we need a model for $P(u|g)$ – a model that tells us what utterances a person would choose conditioned on having a particular goal. LLMs are good at roleplaying Shanahan et al. (2023) and have been demonstrated to possess strong common sense priors Talmor et al. (2022); Zheng et al. (2024). We leverage these strong priors to model $P(u\mid g)$ as an LLM query $\pi_{\text{LLM}}(u,g)$:
@@
Consider maintaining a belief over two possible goals: “I want cocoa” and “I want noodles” given the utterance “I’d like something sweet.” Our method queries an LLM for the likelihood of the utterance based on the prompt “Role-play as a person who wants gg”, where $g$ is replaced with each possible goal. Given a human utterance is “I want a cake”, the likelihood of “I want cocoa” would be higher than “I want noodles”. The sum of the logits for “I want a cocoa” when the LLM is provided that “I want a cake” is the true goal is higher than the sum of the logits for “I want a cake” when the LLM is provided that “I want noodles” is the true goal. See Figure 3 for LLM query details for the inference module (the details about the Unspecified Goal is explained further in the next subsection about the Goal Management Module).
Figure 3: Inference Module LLM Prompts for calculating the log likelihood for each goal by using LLM logits that correspond to the human utterance given context.
Goal Management Module
@@
To implement this module, we need to be able to propose new goals and remove unlikely goals. To propose new goals, a LLM is prompted to return a list of goals that can be added given the current goal list, the conversation transcript, and the robot’s task. We remove unlikely and unsafe goals through two ways. 1) A LLM is prompted to return a list of goals to remove that should not be taken given the robot’s task and the conversation transcript. 2) If a goal is the least likely from our inference methods for more than $n$ rounds of conversation, then it should be removed from the goal list. The goal proposal prompt and the goal removal prompt for the LLM are shown in Figure 4.
Figure 4: LLM prompts involved in the Goal Management Module for proposition of goals and removal of goals that are unlikely or not relevant.
@@
The action module is responsible for two things: 1) selecting action(s) that are available and taking them given the likely goals from the inference module, and 2) checking if the task is completed. This module is domain dependent, and can be easily adapted to other domains and planners. Our implementation for the Grocery Shopper and Robot domains will be further explained in more detail in the next chapter.
@@
Do Action: The primary purpose of this module is for the robot/agent to take actions. We do this by prompting a language model to pick actions given the most likely goals and information about their likelihood magnitudes. If the most likely goal is “Unspecified Goal”, then none of the non-unspecified goals present accurately represent the human preference, so proceed to do “no action”. If another goal is the most likely, then the current goal list represents the human preference well, so pick the top $k$ goals from the goal list. This is a benefit of our method because the action module can take actions that are useful for multiple potential likely goals. The module maintains a record of the actions that have been taken to provide context for other modules. The LLM prompt for Do Action includes information about the previous transcript, the list of possible actions, the most likely goal, and the next $k$ likely goals in order. The prompt will return the list of actions that should be taken in order to accomplish the goal. How the list of possible actions is constructed is explained more in the Experiments section.
@@
Task Completion Check: After a sequence of actions is taken, we need to check if the task is completed according to the human’s satisfaction indicated by their utterance this round or if a certain end action is taken. This completion check is domain and task dependent, and is defined in the Conversation Module. If the task is not completed, then whole pipeline is repeated for another round of conversation.
Figure 5: Method Overview Diagram. The Conversation Module produces single round of dialog. Inference Module Round 1 for determining whether Goal Management Module should add more goals. Unreasonable goals are removed from the goal list. Another round of inference is done to determine the most likely goal(s) to be passed to the action module, which generates an action list for the agent to take. If the task is completed, then the conversation stops.
4 Experiments